Joe is reviewing the investment performance of four mutual funds (A, B, C, and D). Based on a risk-free rate of 3 percent and the systematic risk of each of the portfolios, which one of the funds has the highest systematic risk-adjusted return? (Alternatively, which one has the highest Treynor ratio?)

Return

Std Deviation

Beta

Portfolio A

10.0%

19.0%

0.90

Portfolio B

12.0%

40.0%

0.70

Portfolio C

18.0%

30.0%

1.25

Portfolio D

22.0%

25.0%

1.80
Answer #1

A

Treynor ratio
=(Return-risk free rate)/Beta
=(10-3)/0.9
=7.78

B

Treynor ratio
=(Return-risk free rate)/Beta
=(12-3)/0.7
=12.86

C

Treynor ratio
=(Return-risk free rate)/Beta
=(18-3)/1.25
=12

D

Treynor ratio
=(Return-risk free rate)/Beta
=(22-3)/1.8
=10.56

Portfolio B has the highest value